# RL2Grid: Benchmarking Reinforcement Learning in Power Grid Operations

### v1

@@ 3 RL2Grid Benchmark
State space. Agents have access to the state of the power grid at each time step. The state includes common grid features such as production at each generator; load demands; status, capacity, and cooldown of transmission lines; and the current step. Additional features are provided based on the environment’s characteristics (e.g., maintenance, opponent events, batteries) and the action space. For example, in the topological case, the state includes the topological vector, the connection status of lines, overflow status, and substation cooldowns. In the continuous case, the state includes target and actual dispatches, curtailment, and generator ramping limits. An exhaustive list and description of the features that comprise the state is discussed in Suppl. D.
Reward. The reward is designed to encourage the agent to keep the grid operational for as long as possible while minimizing the (i) overloads and disconnections of transmission lines and (ii) economic costs related to redispatching. To promote good grid operations, the agent gets a cumulative positive constant $R_{\text{survive}}$ for each step, normalized by the total length of a training episode ($\in[0,1]$). The overload reward $R_{\text{overload}}$ is the (normalized) sum of the difference between the flows on the transmission lines and their physical capacity (which goes negative in case of overloads) and is set to a fixed penalty value for disconnected lines in the unconstrained settings ($\in[-1,1]$). The economic costs component $R_{\text{cost}}$ assigns an economic cost to redispatching actions and penalizes energy losses ($\in[-1,0]$). The total reward $R$ an agent gets at each step is then $R=\alpha R_{\text{survive}}+\beta R_{\text{overload}}+\eta R_{\text{cost}}$; the weights are evaluated in Suppl. F.
3.1 Heuristic-guided RL for Expert Grid Operations
Figure 4: Recovery heuristic based on previous L2RPN solutions and operators’ expertise. When a line surpasses the capacity threshold, the RL agent picks actions to address the risky situations. Otherwise, the heuristic computes the recovery actions to revert the grid topology to its original one. We operate on a maximum of one substation per each step as in real operations.
Given the complexity of topological operations, we introduce two heuristics to assist RL agents. These heuristics incorporate operators’ expert actions—topology recovery and idle actions—into the RL training loop, enhancing performance and sample efficiency in complex, high-dimensional grids. While heuristic-guided approaches have been employed by L2RPN competitors (Donnot,, 2020; Marot et al., 2020a, ), no widely adopted solution currently exists that: (i) is easily accessible to RL practitioners, as most solutions are embedded within L2RPN submissions and design choices; and (ii) has been systematically benchmarked across increasingly complex grids to establish a standard for comparison. RL2Grid’s heuristics fill these gaps by integrating operators’ expertise with best practices from L2RPN solutions (e.g., effective recovery actions). This not only offers a strong baseline for comparison with prior work (Marot et al., 2020a, ; Marot et al.,, 2021), but also incorporates valuable insights from human operators, ensuring practical relevance and improved training efficiency.
The first “idle” heuristic (I) does not perform any operation when line capacities $\rho$ are below the safety threshold of $\rho_{\text{max}}=95\%$. The second heuristic is for topology “recovery” (R) (Figure 4), as it restores the grid’s original topology or performs idle actions when line capacities are below the safety threshold. In more detail, when line capacities exceed the safety threshold, the RL agent selects and executes an action based on the current state to bring the grid back to normal operation. When the grid operates normally (i.e., all line flows are under the safety threshold), the recovery policy takes over. If the grid is in its original topology, the heuristic performs an idle action to proceed with the simulation without changes. Otherwise, the heuristic calculates the actions needed to revert each substation to its original configuration. Considering realistic operational limits (i.e., change at most one substation per step), the heuristic first recovers the substation that is most different from its original configuration. Importantly, the recovery actions do not disrupt RL agent training; instead, they emulate typical human operator behavior. Under normal conditions, system operators prefer to maintain the starting topology. However, it is challenging for an RL agent to learn to restore the original topology in high-dimensional action spaces. For this reason, we anticipate that RL baselines augmented with the recovery policy will drastically improve performance and sample efficiency.
3.2 Safe Operations via Constraints
Modeling safe grid operations through constraints is essential, yet prior approaches such as those developed for the L2RPN challenge have not explored CMDP-based tasks for topology control (Marot et al., 2020a, ; Marot et al.,, 2021; Marot et al., 2022b, ; Chen et al.,, 2022; Kelly et al.,, 2020; Chauhan et al.,, 2023; Zhang et al.,, 2020). RL2Grid bridges this gap by introducing CMDP-based environments that reflect critical operational constraints faced by power system operators, ensuring RL agents are trained under realistic, safety-critical conditions. We define two sets of constraints, based on major primary failure modes that can occur in grid operations:
  • Load shedding and islanding constraint (LSI): Failures to meet demand or the creation of electrical islands are treated as a constraint violation, returning a positive cost—this encourages agents to maintain supply-demand balance and ensure grid connectivity consistently. While these are instantaneous constraints in real-world operations, we model them as cumulative constraints (with a zero threshold) to accommodate popular safe RL methods (Su et al.,, 2024; Liu et al., 2021a, ).
    In more detail, let $P_{D,t}$ denote the total demand at time $t$, and $P_{G,t}$ the total generated power. We define the load shedding indicator at $t$ as $L_{t}=\mathbb{1}(P_{G,t}<P_{D,t})$, where $\mathbb{1}(\cdot)$ is the indicator function, which is 1 if the condition inside holds and 0 otherwise. Let $N_{I,t}$ be the number of electrical islands detected in the grid at time $t$ (excluding the main connected component). Define the islanding indicator as $I_{t}=\mathbb{1}(N_{I,t}>0)$. The cost function for LSI at time $t$ is $C_{\text{LSI}}(t)=L_{t}+I_{t}$. Since an episode terminates upon not satisfying the total demand or when an electrical island is detected, we model this instantaneous constraint by setting the constraint threshold to 0 (Stooke et al.,, 2020; Liu et al., 2021b, ; Liu et al., 2021a, ; Su et al.,, 2024; Gu et al.,, 2024). Hence, the policy is considered safe if $C_{\text{LSI}}=0$, meaning no violations occur.
  • Transmission line overload constraint (TLO): Overloading lines risks equipment damage and premature failures. It is therefore important to limit the extent to which overloads occur, which is modeled naturally as a cumulative constraint with a specified threshold. Costs are triggered and accumulate each time a line is overloaded, incentivizing agents to manage power flows effectively and avoid prolonged overloads. Specifically, let $P_{F,\ell,t}$ be the power flow on transmission line $\ell$ at time $t$, and let $P_{F,\ell}^{\max}$ be the thermal limit of line $\ell$. Define the overload indicator for line $\ell$ as $O_{\ell,t}=\mathbb{1}(P_{F,\ell,t}>P_{F,\ell}^{\max})$. If a line is overloaded for more than a certain number of time steps (in our environments, 2 time steps), then that line is disconnected by the environment in order to avoid further damage. Let $D_{\ell,t}=\mathbb{1}(\ell\text{ is disconnected})$ be an indicator function for whether line $\ell$ is disconnected by the environment to avoid damage (this does not include disconnections for scheduled maintenance or that are initiated by the opponent). The total overload cost at time $t$ across all transmission lines $\mathcal{L}$ is then $C_{\text{TLO}}(t)=\sum_{\ell\in\mathcal{L}}(O_{\ell,t}+D_{\ell,t})$. The cumulative constraint formulation over an episode is $\sum_{t=0}^{T}C_{\text{TLO}}(t)<\tau$, where $\tau$ is the (hard-coded) constrained threshold.
These constraints are primarily considered across the 32 topological environments, resulting in additional 64 constrained tasks. [^8] Although RL2Grid allows the use of these constraints in the redispatching and curtailment tasks, we primarily benchmark the topological case, due to the higher relevance to system operators of leveraging RL for topology tasks, and given the large total number of tasks we introduce. By formalizing these operational constraints, RL2Grid aims to foster research on RL methods that balance operational effectiveness with strict safety requirements, establishing, to our knowledge, the first benchmark for safe RL in power grid operations.
@@ 4 Experiments
We assess the performance of popular RL methods that usually serve as building blocks for more complex algorithms in representative tasks of RL2Grid. In particular, we test: (i) DQN (Mnih et al.,, 2013), PPO (Schulman et al.,, 2017), and SAC (Haarnoja et al.,, 2018) (and their heuristic-integrated versions) on the discrete topological action space for the bus14, bus36-MO-v0, bus118-M, bus118-MOB-v0 tasks over all levels of difficulty (i.e., with increasingly bigger action spaces); (ii) PPO, SAC, and TD3 (Fujimoto et al.,, 2018) in the continuous redispatching action space of these environments; (iii) LagrPPO (Stooke et al.,, 2020) in the two constrained versions (LSI and TLO, described in the previous section) of the discrete topological space for the bus14 task. Our experiments address the following key questions: (i) Can commonly-used model-free RL methods deal with high-dimensional power network operations? (ii) What is the impact of integrating existing task-level knowledge as a heuristic-guided policy within these real-world tasks? (iii) How difficult is it to consider constraints when training an RL agent for power grid operations?
Implementation Details. Data collection is performed on Xeon E5-2650 CPU nodes with 64GB of RAM, using CleanRL-based implementations for the baselines (Huang et al.,, 2022). Complete hyperparameters are in Suppl. F. Additionally, we set a strict time limit on the nodes used for data collection, set to 48 hours. As such, different algorithms have different computational requirements, and some of the baselines run for more steps than others. For example, the heuristic-guided methods are more computationally demanding than the baselines given that they often have to check and compute the reversion actions for the power network, and thus run for considerably fewer steps than the baselines. Given the computational resources used, Suppl. E addresses the associated environmental impact of our experiments and our efforts to offset estimated CO2 emissions.
Results.
Table 2: Average survival rate (higher is better) for vanilla RL algorithms (V) and their heuristic versions (I for the idle heuristic, R for the heuristic that restores the grid to its original topology) in representative environments (difficulty 0) with topological actions for DQN, PPO, SAC.
DQN PPO SAC
Env. Diff. V I R V I R V I R
bus14 0 0.07 0.86 0.74 0.74 0.99 0.97 0.17 0.56 0.16
bus36-MO-v0 0 0.04 0.14 0.19 0.06 0.17 0.29 0.01 0.10 0.13
bus118-M 0 0.06 0.17 0.18 0.07 0.13 0.18 0.15 0.18 0.19
bus118-MOB-v0 0 0.07 0.19 0.27 0.04 0.18 0.28 0.01 0.15 0.19
Table 3: Average survival rate (higher is better) of the grid obtained by vanilla RL algorithms (V) in representative environments with continuous redispatching actions for PPO, SAC, TD3.
PPO SAC TD3
Env. Diff. V V V
bus14 0 0.17 0.01 0.06
bus36-MO-v0 0 0.08 0.02 0.01
bus118-M 0 0.18 0.01 0.01
bus118-MOB-v0 0 0.25 0.08 0.07
Due to the high number of tasks in RL2Grid, we report representative results on the bus14 grid at difficulty 0, referring to Suppl. G for the complete results and training curves of the other levels. In particular, the figures show the average survival smoothed over the last 500 episodes of 5 runs per method, with shaded regions representing the standard error; tables show the average survival at convergence (as described above, survival rate denotes the normalized number of time steps the grid remains operational over an episode, with a survival rate of 1 indicating one month of successful grid operations). Table 2 shows the preliminary results of our topological action spaces evaluation. We indicate with V the original (“vanilla”) model-free baseline, and with I and R the baseline augmented with the heuristics described in Section 3.1 (idle and reconnect, respectively). For these topological cases, we report the results for the first level of difficulty (i.e., level 0, considering 50 discrete actions), and Table 3 shows the results for the redispatching action spaces. [^9] We recall the redispatching case only has one level of difficulty since it only considers one action for each generator. Overall, we notice these popular RL baselines struggle to deal with the complexities of power network operations described in Section 2. Considering the lower number of training steps, we also notice that the heuristic-guided versions of the baselines typically achieve higher performance, despite being not nearly sufficient to operate the grid for long periods of time in complex grids.
Figure 5 compares the training performance for model-free baselines, their heuristic-enhanced versions, and constrained variants in the representative bus14 task, which involves topological actions at difficulty level 0. Without the aid of constraints and heuristics, PPO is the only algorithm that successfully learns a good policy for this relatively small grid. However, incorporating human-informed heuristic operations significantly improves performance and sample efficiency across all algorithms. With heuristic augmentation, PPO is able to operate the grid successfully over extended periods, while DQN and SAC also demonstrate strong overall performance. Interestingly, our results suggest that performing idle operations instead of reverting the grid to its original topology increases performance for RL algorithms. On the other hand, enforcing instantaneous or cumulative constraints makes the task considerably more difficult, as agents are penalized for any violations. As a result, LagrPPO struggles even on the bus14 system, failing to achieve effective control while still frequently exceeding the imposed constraint thresholds (more details in Suppl. G).
These results motivate the need for further advancements in RL algorithms that can contend with the complex dynamics and aleatoric uncertainty, long-horizon goals, and hard physical constraints represented within these tasks. By providing common ground for the community, we hope to foster further research on these fronts.
@@ Results.
Figure 5: Average survival comparison of model-free RL baselines (left), heuristic-augmented versions (center), and constrained variants (right) on the bus14 task with topological actions at difficulty level 0 (higher values are better). Incorporating heuristics significantly enhances performance and sample efficiency. In contrast, enforcing operational constraints introduces additional challenges, leading to poor performance while failing to satisfy the imposed safety thresholds.
Performance analysis. Figure 6 analyzes how the agent is learning to control the grid. In particular, for each setup considered in the previous section (represented in each row), we show the following three operational components—for clarity, these are visualized as additional reward signals:
  • Margin (first column): Represents the cumulative available margin across all transmission lines, where disconnections are penalized and lower power usage is rewarded. Higher values indicate that the agent maintains greater flexibility to handle contingencies. Overall, we find that successful agents tend to maximize line margins.
  • Overloads (second column): Measures a cumulative penalty based on the severity and duration of transmission line overloads. Lower values imply that the agent maintains power flows within safe operational limits, which is a key characteristic of effective policies. Unsurprisingly, the unconstrained agents violate the overload constraints. Perhaps surprisingly, the constrained versions actually exhibit even more severe violations of these constraints, which we hypothesize is due to their overall difficulty in optimizing performance in the constrained environment.
  • Topology (third column): Quantifies deviations from the initial grid configuration (where all elements are connected to the first bus). A lower value suggests minimal changes, while higher values indicate significant topological modifications. We find that successful agents tend to actively reconfigure the grid to optimize operations.
@@
Figure 6: Average (non normalized) reward for key operational components across model-free, heuristic-based, and constrained setups (rows) during training. The first column represents line margins, where higher values indicate better contingency management. The second column tracks overload penalties, with lower values reflecting improved grid stability. The third column captures topology modifications, showing the extent to which agents reconfigure the grid.

### v2

@@ 3 RL2Grid Benchmark
State space. Agents have access to the state of the power grid at each time step. The state includes common grid features such as production at each generator, load demands, status, capacity, and cooldown of transmission lines, and the current step. Additional features are provided based on the environment’s characteristics (e.g., maintenance, opponent events, and/or batteries—see Table 1) and the action space. In the topological case, the state includes the topological vector (an integer vector indicating where each device is connected), the connection status of lines, overflow status, and substation cooldowns. In the continuous case, the state consists of target and actual dispatches, curtailment, and generator ramping limits. These are the main features used by the AC power flow solver to transition the grid to the next state. Due to space limitations, an exhaustive list and description of the features that comprise the state is discussed in Appendix D.
Action spaces. Each grid has two types of tasks, depending on the nature of their action space.
(i) Topology space: Agents take discrete actions that modify the topology of the substations—disconnecting or reconnecting a line, or changing the bus to which a component is connected. [^4] Human operators currently modify the grid topology manually based on historical behaviors; there is no tractable approach to obtain optimal topology optimization solutions (at scale) as of yet. Line switching introduces one discrete action per line, whereas bus reassignments (or “bus-splitting”) yield an exponentially large number of valid actions depending on the number of elements connected to the substations. Specifically, the topological action space for a double bus substation composed of $N_{\text{lines}}$ lines, $N_{g}$ generators, $N_{l}$ loads, has size $N=2^{N_{\text{lines}}+N_{g}+N_{l}-1}-1$ (Chauhan et al.,, 2023). For instance, substation #5 in Figure 2 has 7 elements, resulting in 63 possible actions, while in the larger bus36 and bus118 grids, a single substation can have over 65,000 possible configurations.
Figure 3: Action ranking for the bus14 task. The dashed line separates the difficulty levels (i.e., with 50 and 209 actions).
Considering the size of the space, we propose “difficulty levels” in which the action space has an increasing number of topology actions. We selected these action spaces through extensive simulations (72 hours on the computer cluster detailed in Section 4) by ranking the full action space based on the survival rate for the grid. This rate represents the number of steps that each action maintains the grid in normal conditions over an episode, and is defined as the normalized number of steps for which an action does not cause a grid collapse (because the total demand is not satisfied or parts of the grid become disconnected). In detail, we uniformly sampled topological actions, and after ordering them by highest survival rate, we took the first $N_{\text{actions}}$ from the ordered space, where $N_{\text{actions}}$ increases with each difficulty level. For example, Figure 3 shows the ranking for the bus14 grid, highlighting how suitable all the topological actions are to address grid contingencies (and how easier levels contain actions that are more likely to address a contingency). To further motivate our method, we visually analyze the impact of the resultant action spaces in Appendix C, where we also summarize the difficulty levels and the size of their discrete action spaces. Considering these levels, RL2Grid has a total of 32 topology-based tasks.
(ii) Redispatching and curtailment space: Agents take continuous actions changing how power generation is scheduled. Unlike topology actions, which only involve remotely activating a switch, these actions are not free. Economic costs arise from altering the planned generation schedule of power plants, increased fuel costs, and financial compensation for renewable energy producers. Redispatching actions apply to fossil fuel-based generators, while curtailment actions apply to renewable energy-based generators. Batteries, if present, are also considered generators and come with continuous actions for charging/discharging operations. This action space is relatively tractable for RL algorithms since it involves one continuous action per generator (i.e., $N=N_{g}$). Thus, we present a total of 7 continuous action-based training environments (one per base grid).
Reward. The reward design is informed by TSOs’ mandate to satisfy real-world grid operation requirements: promote long-term safety and efficiency by rewarding grid survival and penalizing unsafe or costly actions (in terms of economic costs). At each step $t$, the reward an agent gets is $R_{t}=\alpha R_{\text{survive},t}+\beta R_{\text{overload},t}+\eta R_{\text{cost},t}$, where the weights are evaluated in Appendix F. Specifically, the agent gets a cumulative positive constant $R_{\text{survive},t}$ for each step, normalized by the total length of a training episode ($\in[0,1]$). The overload and cost rewards are defined as:
(i) Overload: Penalizes line overloads and disconnections, and rewards available line capacity based on the difference between line flows and capacity limits. In unconstrained settings, disconnected lines incur a fixed penalty. This is more formally defined as:
$R_{\text{overload},t}=\sum_{\ell\in\mathcal{L}}\left[\max\left(0,\frac{P_{F,\ell,t}-P_{F,\ell}^{\max}}{P_{F,\ell}^{\max}+\epsilon}\right)-\mathbbm{1}(\ell\text{ is disconnected})\right],$
where $P_{F,\ell,t}$ is the power flow on line $\ell$ at time $t$, $P_{F,\ell}^{\max}$ is its capacity limit, $\epsilon$ is a small constant to avoid divisions by 0, and the indicator function returns 1 if the line is disconnected. This term is then normalized to lie within $[-1,1]$.
(ii) Cost: Penalizes redispatching or curtailment actions based on deviations from planned dispatch schedules and energy losses. This is defined as:
$R_{\text{cost},t}=-\left[\left(P_{G,t}-P_{D,t}\right)+|c_{\text{redisp},t}|+|P_{\text{storage},t}|\right]c_{\text{marginal},t},$
where $P_{G,t}$ and $P_{D,t}$ denote the total power generated and total demand consumed at time $t$, respectively, with their difference representing transmission losses, $c_{\text{redisp},t}$ corresponds to the redispatched power (i.e., the absolute deviation from scheduled generator setpoints), and $P_{\text{storage},t}$ represents the power exchanged with storage units. All cost components are scaled by the marginal generation cost $c_{\text{marginal},t}$, defined as the cost per MWh of the most expensive generator currently producing power. This value is also normalized to lie in the range $[-1,0]$.
3.1 Heuristic-guided Transitions
Figure 4: Recovery heuristic combining L2RPN strategies and operator expertise. The agent acts under risk (e.g., line overloads); otherwise, the heuristic incrementally restores the original topology.
To reduce the problem horizon and (potentially) improve learning stability and sample efficiency, RL2Grid incorporates two expert-informed heuristics that modify the transition dynamics described above. These heuristics emulate human operator behavior: they suppress agent actions when the grid is stable and allow the agent to try to recover normal operations when contingencies occur.
The idle heuristic (I) is triggered when all line loadings are below a safety threshold set to $95\%$ of their capacity. In this case, the agent’s control is suspended and replaced by no-op actions. The recovery heuristic (R), shown in Figure 4, activates when the grid is safe but the topology differs from its original configuration. It incrementally restores the original topology, modifying at most one substation per step, and defaults to idle once recovery is complete.
These heuristics modify the environment so that each agent action initiates a sequence of $n$-step heuristic-guided transitions during which rewards accumulate. While our heuristics encode operational priors from winning L2RPN competitors (Marot et al., 2020a, ; Marot et al.,, 2021), our methods have the benefit of being embedded directly in the environment for compatibility with standard RL libraries and are systematically benchmarked. We believe this design balances expert-in-the-loop guidance with the learning process, enabling more sample-efficient training and improved policy performance in realistic, safety-critical scenarios.
3.2 Fostering Safe Operations via Constrained RL
While prior works such as L2RPN have not incorporated constrained formulations to foster safety (Marot et al., 2020a, ; Marot et al.,, 2021; Marot et al., 2022b, ), RL2Grid introduces CMDP-based tasks that reflect two key classes of safety violations faced by system operators as constraints.
In the load shedding and islanding (LSI) case, unsatisfied demand and islanding trigger a positive cost. Let us denote the total demand and generation at time $t$ as $P_{D,t}$ and $P_{G,t}$, respectively, and define $L_{t}=\mathbbm{1}(P_{G,t}<P_{D,t}),\quad I_{t}=\mathbbm{1}(N_{I,t}>0),$ where $N_{I,t}$ is the number of disconnected components. The LSI cost is $C_{\text{LSI}}(t)=L_{t}+I_{t}$, with a zero cumulative threshold $\sum_{t}C_{\text{LSI}}(t)=0$ to model a hard safety constraint. For transmission line overload (TLO), a positive cost occurs upon thermal overloads and line disconnections. Let us denote the power flow on line $\ell$ as $P_{F,\ell,t}$, its capacity as $P_{F,\ell}^{\max}$, and define $O_{\ell,t}=\mathbbm{1}(P_{F,\ell,t}>P_{F,\ell}^{\max}),\quad D_{\ell,t}=\mathbbm{1}(\ell\text{ is disconnected})$, where $D_{\ell,t}$ excludes scheduled maintenance or opponent-driven disconnections. The TLO cost is $C_{\text{TLO}}(t)=\sum_{\ell\in\mathcal{L}}(O_{\ell,t}+D_{\ell,t})$, with a cumulative threshold $\sum_{t=0}^{T}C_{\text{TLO}}(t)<\tau$ to model this as a “soft” constraint.
These constraints are applied across all 32 topology-based environments, resulting in 64 additional constrained variants. [^5] Constraints are compatible with redispatching tasks but are primarily evaluated on topology due to their operational relevance and the complexity of associated actions. By modeling these critical safety violations, RL2Grid enables the development and benchmarking of safe RL methods for real-world grid operations.
@@ 4 Experiments
We evaluate the performance of baseline RL algorithms that typically serve as building blocks for more complex algorithms in representative RL2Grid tasks. In particular, we test: (i) (double) DQN (van Hasselt et al.,, 2016), PPO (Schulman et al.,, 2017), and SAC (Haarnoja et al.,, 2018) and their heuristic versions on the discrete topological action space for the bus14, bus36-MO-v0, bus118-M, bus118-MOB-v0 tasks over most levels of difficulty; (ii) PPO, SAC, and TD3 (Fujimoto et al.,, 2018) in the continuous redispatching action space of these environments; (iii) and the Lagrangian version of PPO, LagrPPO (Stooke et al.,, 2020), in the two constrained versions (LSI and TLO) of the topological bus14 task. We consider these tasks to be representative, as they provide sufficient empirical evidence of the current performance of the baselines in power grid operations.
Implementation and Data Collection. Data collection is performed on Xeon E5-2650 CPU nodes with 256GB of RAM, using CleanRL-based implementations for the baselines (Huang et al.,, 2022) and the hyperparameters—selected via grid search—in Appendix F. If not specified otherwise, the results show the average survival smoothed over 500 episodes of 10 runs per method, with shaded regions representing the $95\%$ confidence intervals. As described above, the survival denotes the normalized number of time steps the grid remains operational over an episode, with a survival rate of 1 indicating one month of successful grid operations. We set a strict time limit on the nodes used for data collection, set to 48 hours for each individual run. The experiments in this work (excluding the hyperparameter search) required a total of $>$180,000 CPU hours to execute, and Appendix E addresses the associated environmental impact and our efforts to offset estimated CO2 emissions.
@@
Figure 5: Average survival of the baselines (left), heuristic versions (center), and constrained variants (right) on the bus14 task with topological actions at difficulty 0 (higher values are better).
Table 2: Average survival rate (higher is better) in a subset of difficulty 0 tasks with topological actions for the baselines: idle (I), MILP optimization (MILP); and RL algorithms: vanilla RL (V), and the heuristic versions (I, R) for DQN, PPO, and SAC.
Baseline DQN PPO SAC
Env. Diff. I MILP V I R V I R V I R
bus14 0 0.18 0.06 0.07 0.86 0.74 0.74 0.99 0.97 0.17 0.56 0.16
bus36-MO-v0 0 0.03 0.06 0.04 0.14 0.19 0.06 0.17 0.29 0.01 0.10 0.13
bus118-MOB-v0 0 0.10 0.05 0.07 0.19 0.27 0.04 0.18 0.28 0.01 0.15 0.19
Table 3: Average survival rate (higher is better) of the grid in a subset of tasks with continuous redispatching actions obtained by an idle baseline (I) and RL-based vanilla baselines (V) for PPO, SAC, and TD3.
Env. Diff. I PPO (V) SAC (V) TD3 (V)
bus14 0 0.00 0.17 0.01 0.06
bus36-MO-v0 0 0.08 0.08 0.02 0.01
bus118-MOB-v0 0 0.11 0.25 0.08 0.07
Results. We indicate with V the “vanilla” baselines, and with I and R the experiments with the heuristics described in Section 3.1. Overall, the RL baselines struggle to deal with the real-world complexities of power grid operations. As expected, we also notice that the heuristic-guided transitions reduce the problem complexity and typically achieve higher performance, despite being not nearly sufficient to operate complex grid setups for long periods of time.
Figure 5 compares the training performance of the baseline algorithms, the heuristic versions, and the constrained variants on the topological bus14 grid at difficulty level 0. Among the unconstrained baselines, only PPO successfully learns an effective policy in this relatively simple environment. However, incorporating human-informed heuristic operations leads to notable improvements in both performance and sample efficiency across all methods. With heuristic augmentation, PPO achieves good long-term control, while DQN and SAC also exhibit strong performance. Interestingly, our results indicate that performing idle operations in smaller grid domains—rather than reverting to the original topology—can improve RL performance. In contrast, introducing constraints significantly increases task difficulty: agents are penalized for violations, and LagrPPO struggles even in this basic setting, failing to learn effective control and frequently exceeding constraint thresholds (see Appendix G for details). Finally, Tables 2 and 3 report the average survival at convergence for the unconstrained baselines in both topological and redispatching tasks. [^6] Due to space constraints, we report only average performance. Full results and training curves, including the bus118-M task and higher difficulty levels, are available in Appendix G. For topology control, we evaluate two traditional baselines: an “idle” policy (I) and the MILP-based agent from Grid2Op (RTE France,, 2022), which minimizes line overloads via topological actions under DC power flow approximations. [^7] This approximation is necessary due to the intractability of AC formulations (Marot et al., 2020a, ; Marot et al.,, 2021). For redispatching, only the idle baseline is considered, as no built-in agent is provided in Grid2Op. Notably, even model-free RL-based agents consistently outperform these traditional methods across both settings.
These performance results motivate the need for further advancements in RL algorithms that can contend with the complex dynamics and aleatoric uncertainty, long-horizon goals, and hard physical constraints of real-world tasks such as power grid operations.
@@
Figure 6: Average score for key operational components across vanilla baselines. The first column represents line margins, where higher values indicate better contingency management. The second column tracks overload penalties, with lower values reflecting improved grid stability. The third column captures topology modifications, showing the extent to which agents reconfigure the grid.
Performance analysis. Figure 6 analyzes how the baseline algorithms are learning to control the grid. [^8] A similar analysis for the heuristic and constrained experiments is available in Appendix H. In particular, we quantify three distinct operational metrics—maring, overloads, and topology—each defined and shown (with $95\%$ confidence intervals) as a scalar-valued “score" signal derived from the agent’s interaction with the environment:
  • Margin (first column): Represents the cumulative available margin across all transmission lines, where disconnections are penalized and lower power usage is rewarded. It is defined per-step $t$ as: $\sum_{\ell\in\mathcal{L}}m_{\ell,t}$, where $m_{\ell,t}=\max\left(0,P_{F,\ell}^{\max}-|P_{F,\ell,t}|\right)$, if line $\ell$ is connected, $m_{\ell,t}=-1$ if not. Higher values indicate that the agent maintains greater flexibility to handle contingencies. Overall, we find that successful agents tend to maximize line margins.
  • Overloads (second column): Shows the overload component $R_{\text{overload,t}}$ as defined in Section 3. Lower values imply that the agent maintains power flows within safe operational limits, which is a key characteristic of effective policies. Unsurprisingly, the unconstrained agents violate the overload constraints. Perhaps surprisingly, the constrained versions actually exhibit even more severe violations of these constraints, which we hypothesize is due to their overall difficulty in optimizing performance in the constrained environment.
  • Topology (third column): Quantifies deviations from the initial grid configuration (where all elements are connected to the first bus). It is defined per-step $t$ as: $-d(G_{t},G_{0})$, where $G_{t}$ denotes the grid topology at time $t$, and $d(G_{t},G_{0})$ is the Hamming distance of the topology at time $t$ from the initial topology $G_{0}$. This metric quantifies the extent of grid reconfiguration. Negative values indicate more significant structural interventions, which often correlate with better operational performance in learned policies. A value close to 00 suggests minimal changes, while higher values indicate significant topological modifications. We find that successful agents tend to actively reconfigure the grid to optimize operations.